OpenStudy (anonymous):

Solve the equation : 5•3^2x=25

OpenStudy (littlebird):

is that (3^2)x or 3^(2x)?

OpenStudy (anonymous):

3^(2x)

OpenStudy (littlebird):

I'm going to start with this hint then: ln(M^n) = n(ln(M))

OpenStudy (littlebird):

the answer is going to look weird because of the natural log

OpenStudy (anonymous):

So I would put 2x(ln(3)) ?

OpenStudy (littlebird):

5 * 3^(2x) = 25 3^(2x) = 5 ln(3^(2x)) = ln(5) 2xln(3) = ln(5) x = ln(5)/2ln(3)

OpenStudy (anonymous):

Okay thanks